### v2

@@ Appendix A Spin connection in curved space
Given a manifold $M$ equipped with metric $g$, without loss of generality we choose orthogonormal frame $\{e_{i}\}$ and the dual frame $\{\theta^{i}\}$, the connection 1-form is ${\omega_{j}}^{i}$, with which the covariant derivative can be written as
$De_{i}={\omega_{i}}^{j}\otimes e_{j}$ (A.1)
The corresponding connection form $\omega$ on the frame bundle $F(M)$ has following relation with connection 1-form on base manifold $M$
$\sigma^{*}\omega:=\begin{pmatrix}{\omega_{1}}^{1}&{\omega_{2}}^{1}&\cdots&{\omega_{n}}^{1}\\ {\omega_{1}}^{2}&{\omega_{2}}^{2}&\cdots&{\omega_{n}}^{2}\\ \vdots&\vdots&\ddots&\vdots\\ {\omega_{1}}^{n}&{\omega_{2}}^{n}&\cdots&{\omega_{n}}^{n}\end{pmatrix}:T_{p}M\to\mathfrak{o}(n),$ (A.2)
where $\sigma:M\to F(M),p\mapsto\{p,e_{1},e_{2},\cdots,e_{n}\}\in F(M)$. Supposing the first Stiefel-Whitney class $w_{1}(M)$ and the second Stiefel-Whitney class $w_{2}(M)$ are all zero so that the frame bundle $F(M)$ can be reduced to principle bundle $Sp(M)$, whose structure group is spin group $Spin(n)$, about the homomorphism $\rho:Spin(n)\to SO(n)$.
Using the homoporphism $\rho$ we can get the connection form on $Sp(M)$ which is induced from the connection form $\omega$ on $F(M)$. The local connection form $\hat{\omega}:T_{p}M\to\mathfrak{spin}(n)$ of $Sp(M)$ is $(\rho_{*})^{-1}\circ\sigma^{*}\omega$, Levi-Civita connection means the connection 1-forms are anti-symmetry ${\omega_{j}}^{i}=-{\omega_{i}}^{j}$, so that the local connection form can be written as
@@
where $e_{i}e_{j}\in Cl(T_{p}M)$. The spin representation $\tau:Spin(n)\to End(\Psi)$ induces the vector bundle called spin bundle
$S(M):=Sp(M)\times_{\tau}\Psi.$ (A.4)
Recall the definition of Clifford bundle[56], we claim the covariant differentiation of a spinnor field $\Psi$ is
$D\Psi=d\Psi+\frac{1}{2}\sum_{i<j}{\omega_{i}}^{j}e_{i}\cdot e_{j}\cdot\Psi=d\Psi+\frac{1}{4}\sum_{ij}{\omega_{i}}^{j}(e_{i}\cdot e_{j}-e_{j}\cdot e_{i})\cdot\Psi,$ (A.5)
where $\cdot$ is Clifford multiple. Using the gamma matrix we can rewrite it in a more familiar form, i.e. redefine $\Gamma^{a}:=\eta^{ab}e_{b}\cdot$ in Lorentz manifold.
$D\Psi=d\Psi+\frac{1}{4}\sum_{ab}{\omega_{a}}^{b}[\Gamma^{a},\Gamma^{b}]\Psi,$ (A.6)
where $\eta={\rm diag}(1,-1,-1,-1,-1)$ and the inner-product in the spinor space is $(\Psi_{1},\Psi_{2}):=\Psi_{1}^{\dagger}\gamma^{0}\Psi_{2}$, so we can check the proposition of Clifford bundle $\{\Gamma^{a},\Gamma^{b}\}=-2\eta^{ab},~{}(\Psi_{1},\Gamma^{a}\Psi_{2})=-(\Gamma^{a}\Psi_{1},\Psi_{2})$, and the Dirac operator is $\Gamma^{a}D_{e_{a}}$. So that the equation of motion of a free fermion with mass $m\in\mathbb{R}$ is
@@
In particular, if the metric has the easy form $g=\sum_{a}\eta^{aa}g_{aa}dx^{a}dx^{a}$, we can write the connection 1-form explicitly. Following is the orthogonormal frame and its dual frame.
$\theta^{a}=\sqrt{g_{aa}}dx^{a},~{}e_{a}=\frac{1}{\sqrt{g_{aa}}}\frac{\partial}{\partial x^{a}}.$ (A.8)
Then using Cartan structure equation $d\theta^{a}=\theta^{b}\wedge{\omega_{b}}^{a}$, we can write the connection 1-form ${\omega_{b}}^{a}$
@@
${\omega_{b}}^{a}=\frac{1}{\sqrt{g_{bb}}}\frac{\partial\sqrt{g_{aa}}}{\partial x^{b}}dx^{a}-\frac{1}{\sqrt{g_{aa}}}\frac{\partial\sqrt{g_{bb}}}{\partial x^{a}}dx^{b},~{}~{}else$ (A.9)
@@
$D_{e_{c}}\Psi=\frac{1}{\sqrt{g_{cc}}}\frac{\partial\Psi}{\partial x^{c}}+\frac{1}{4}\sum_{ab}\braket{{\omega_{a}}^{b},e_{c}}[\Gamma^{a},\Gamma^{b}]\Psi$ (A.10)
@@
In the article we denote $D_{e_{a}}$ as $D_{a}$ in short.
@@ Appendix B Chern class
Chern class is a class in $H_{dR}^{2k}(M;\mathbb{R})$ which can indicate the structure of complex vector bundle $(E,M,\pi)$, specific details are following[57].
@@
Supposing $V$ is a finite dimension linear space, for $k\geq 1$ let’s consider following map
$f:V\times V\times\cdots\times V\to\mathbb{R},~{}(v_{1},v_{2},\cdots,v_{k})\mapsto f(v_{1},v_{2},\cdots,v_{k})$
If $f$ meets following conditions:
@@
    $f$ is symmetrical, i.e. $f(v_{\sigma(1)},v_{\sigma(2)},\cdots,v_{\sigma(k)}),~{}\forall\sigma\in S_{k}$
All these $f$ compose linear space $S^{k}(V^{*})$, denote $S^{0}(V^{*})=\mathbb{R}$.
$G$ is a Lie group, whose Lie algebra is $\mathfrak{g}$, supposing $f\in S^{k}(\mathfrak{g}^{*})$, if
$f(Ad_{g^{-1}}X_{1},Ad_{g^{-1}}X_{2},\cdots,Ad_{g^{-1}}X_{k})=f(X_{1},X_{2},\cdots,X_{k})$
$f$ is an invariant polynomial.
$I^{k}(G):=\{f\in S^{k}(\mathfrak{g}^{*}):f~{}is~{}invariant~{}polynomial\}$
Denote $I^{*}(G)$ as all the invariant polynomial. Specially, if $\eta_{1},\eta_{2},\cdots,\eta_{k}$ are $V$-value differential forms on a manifold, select a basis $\{E_{i}\}$ of $V$
$\eta_{i}={\eta_{i}}^{j}E_{j}$
$f(\eta_{1},\eta_{2},\cdots,\eta_{k}):=f(E_{j_{1}},E_{j_{2}},\cdots,E_{j_{k}}){\eta_{i}}^{j_{1}}\wedge{\eta_{2}}^{j_{2}}\wedge\cdots\wedge{\eta_{k}}^{j_{k}}$
In the case under discussion, the properties about $I^{k}(GL(r,\mathbb{R}))$ is valuable.
Proposition B.1
Given $f\in I^{k}(GL(r,\mathbb{R}))$, these properties are tenable.
@@
    $\forall X,X_{i}\in\mathfrak{gl}(r,\mathbb{R})$
    $f([X_{1},X],X_{2},\cdots,X_{k})+\cdots+f(X_{1},X_{2},\cdots,[X_{k},X])=0$
@@
    $\eta,\eta_{i}$ are all $\mathfrak{gl}(r,\mathbb{R})$-value differential forms whose order are all odd numbers, then
    $f([\eta_{1},\eta],\eta_{2},\cdots,\eta_{k})+\cdots+f(\eta_{1},\eta_{2},\cdots,[\eta_{k},\eta])=0$
@@
As f is an element in $I^{k}(GL(r,\mathbb{R}))$, for all $t\in\mathbb{R}$
$f(\exp(-tX)X_{1}\exp(tX),\cdots,\exp(-tX)X_{k}\exp(tX))=f(X_{1},X_{2},\cdots X_{k})$
Differentiate with respect to $t$ at $t=0$. The first property is proofed. Using this property and choosing a basis in $\mathfrak{gl}(r,\mathbb{R})$ to expand $\eta,\eta_{i}$, the second one can also be verified.
$(E,M,\pi)$ become as a vector bundle, whose structure group is $GL(r,\mathbb{R})$, is equipped with an arbitrary choice of connection, denoted as $D$,
$De_{i}={\omega_{i}}^{j}\otimes e_{j},$ (B.1)
where $\{e_{i}\}$ is a frame in $E$ and ${\omega_{j}}^{i}$ are connection forms.These connection forms can be rewritten into a matrix
@@
$\omega$ is a $\mathfrak{gl}(r,\mathbb{R})$-value 1-form, with its curvature defined as a $\mathfrak{gl}(r,\mathbb{R})$-value 2-form, which is:
$\Omega:=d\omega+\frac{1}{2}[\omega,\omega]$ (B.3)
Let f is an element in $I^{k}(GL(r,\mathbb{R}))$, then
$f(\Omega):=f(\Omega,\Omega,\cdots,\Omega)$ (B.4)
Considering f is invariable, $f(\Omega)$ is well defined across the entire base manifold $M$. Further investigation is warranted into the properties of $f(\Omega)$, which are listed below.
Proposition B.2
For all connection on $(E,M,\pi)$, $f(\Omega)$ is a $2k$ order close form on $M$. The cohomology class $[f(\Omega)]$ which $f(\Omega)$ represents in $H_{dR}^{2k}(M)$ is independent to the choice of connection.
@@
Check if $f(\Omega)$ is closed, with
$df(\Omega)=f(d\Omega,\Omega,\cdots,\Omega)+f(\Omega,d\Omega,\cdots,\Omega)+\cdots+f(\Omega,\Omega,\cdots,d\Omega)$
and Bianchi identity, we have
$df(\Omega)=f([\Omega,\omega],\Omega,\cdots,\Omega)+f(\Omega,[\Omega,\omega],\cdots,\Omega)+\cdots+f(\Omega,\Omega,\cdots,[\Omega,\omega])=0$
So $f(\Omega)$ is truly a closed form. Then given two connection forms $\omega_{0},\omega_{1}$ on M, define $\omega_{t}:=(1-t)\omega_{0}+t\omega_{1}$, we need to demonstrate $f(\Omega_{1})-f(\Omega_{0})=d\Phi$ to ensure their difference is a exact form.
@@
Where $\eta:=\omega_{1}-\omega_{0}$, then using proposition 2.2 and Bianchi identity, we can get the proof.
@@
Denote $w(E,f)=[f(\Omega)]\in H_{dR}^{2k}(M)$, the map $w(E,\cdot):I^{*}(G)\to H_{dR}^{*}(M)$ is called Chern-Weil homomorphism. Subsequently, the Chern characteristic class can be defined.
Definition B.2 (Chern class)
$U(n)$ is a subgroup of $GL(n,\mathbb{C})$, and its Lie algebra is $\mathfrak{u}(n)$,
$U(n):=\{g\in GL(n,\mathbb{C}:gg^{\dagger}=I_{n}\},~{}\mathfrak{u}(n):=\{A\in\mathfrak{gl}(n,\mathbb{C}):A=-A^{\dagger}\}$
Investigate following expression
$\det\left(\lambda I_{n}-\frac{1}{2\pi i}A\right)=\sum_{k=0}^{n}c_{k}(A)\lambda^{n-k},~{}\lambda\in\mathbb{C}$
$c_{k}(mA)=m^{k}c_{k}(A)$, the determinant doesn’t change under unitary transformation, i.e. $c_{k}$ is Ad invariant, it is also easy to check $c_{k}$ is real. so $c_{k}$ can determine a k order polynomial of A, we still call it $c_{k}$.
$c_{k}(v_{1},v_{2},\cdots,v_{k}):=\frac{1}{k!}\frac{\partial^{k}}{\partial t_{1}\partial t_{2}\cdots\partial t_{k}}c_{k}(t_{1}v_{1}+t_{2}v_{2}+\cdots+t_{k}v_{k}),~{}t_{i}\in\mathbb{R},~{}v_{i}\in\mathfrak{u}(n)$
Denote $c_{k}(E)=w(E,c_{k})=[c_{k}(\Omega)]\in H_{dR}^{2k}(M;\mathbb{R})$, we call $c_{k}(E)$ is the k-th Chern characteristic class, it isn’t related to the choice of connection.
Appendix C Euler characteristic number and Chern number in 2-dimensional system
After given the definition of Chern class, we will give an example in 2-dimensional system, where Chern number is closely related to Euler’s characteristic number. Through following discussion it will be more reliable that Chern class truly reflects the topological properties of the system.
For a specific d-dimensional system with the Hamiltonian $\hat{H}(k),~{}k\in\mathbb{T}^{d}$, consider one branch of the eigen values $E_{n}(k)$ and corresponding eigen vectors |$\ket{n(k)}$, which satisfy the eigen equation $\hat{H}(k)\ket{n(k)}=E_{n}(k)\ket{n(k)}$ for any $k$ in Brillouin zone $\mathbb{T}^{d}$. Notice that the eigen vectors permit a phase transformation, i.e. for all $\alpha\in\mathbb{C}$ the $\alpha\ket{n(k)}$ also satisfy the eigen equation, so we can define a complex vector bundle $V_{n}$ on the Brillouin zone as
$V_{n}:=\bigcup_{k\in\mathbb{T}^{d}}\{\alpha\ket{n(k)}|\alpha\in\mathbb{C},~{}\hat{H}(k)\ket{n(k)}=E_{n}(k)\ket{n(k)}\}.$ (C.1)
Then we can consider the Chern class of $V_{n}$. Obviously, the structure group for $V_{n}$ is $U(1)$, then using the definition of Chern class, the $c_{1}$ Chern class for $V_{n}$ is
$\det\left(\lambda-\frac{A}{2\pi i}\right)$ $=$ $c_{0}(A)\lambda+c_{1}(A),~{}A\in\mathfrak{u}(1),$
$c_{1}(V_{n})$ $=$ $-\frac{[\Omega]}{2\pi i}.$ (C.2)
Where $\Omega$ is the curvature for $V_{n}$. As a result, we can get the Chern class for vector bundle $V_{n}$ as long as we have a connection on $V_{n}$. Without loss of generality we choose the connection form $\omega$ and covariant differentiation $D$ as following
$\omega:=i\bra{n(k)}d\ket{n(k)}:T_{k}\mathbb{T}^{d}\to\mathfrak{u}(1),~{}D\ket{n(k)}=\omega\otimes\ket{n(k)}.$ (C.3)
It can be easily checked that this connection satisfy the definition of connection and using Cartan’s structure equation to get corresponding curvature form $\Omega$ as
$\Omega=d\omega-\omega\wedge\omega=id\bra{n(k)}\wedge d\ket{n(k)},$ (C.4)
and the corresponding $c_{1}$ Chern class with the condition that all eigen vectors has been normalized
$c_{1}=-\frac{[\Omega]}{2\pi i}=-\frac{[d\bra{n(k)}\wedge d\ket{n(k)}]}{2\pi}.$ (C.5)
As it has been discussed, the $c_{1}$ class is not affected by the choice of connection. Then if $d=2$, i.e. this is a 2-dimensional system, we can integrate the $c_{1}$ Chern class all over the Brillouin zone to get the Chern number $C(E_{n})$ of the band $E_{n}$
$C(E_{n}):=\int_{\mathbb{T}^{2}}c_{1}=-\int_{\mathbb{T}^{2}}\frac{d\bra{n(k)}\wedge d\ket{n(k)}}{2\pi}.$ (C.6)
Here we integrate the curvature form directly, for the integration of any exact form is 00.
In another perspective, the eigen subspace $\{\alpha\ket{n(k)}|\alpha\in\mathbb{C},~{}\hat{H}(k)\ket{n(k)}=E_{n}(k)\ket{n(k)}\}$ corresponding to energy band $E_{n}$ is a complex linear space, we can realify the complex linear space to be a real linear space.

### v3

@@ Appendix A Spin connection in curved space
Given a manifold $M$ equipped with a metric $g$, we may choose an orthonormal frame $\{e_{i}\}$ and its dual coframe $\{\theta^{i}\}$ without loss of generality. The connection 1-form ${\omega_{j}}^{i}$ defines the covariant derivative as
$De_{i}={\omega_{i}}^{j}\otimes e_{j}.$ (A.1)
The corresponding connection form $\omega$ on the frame bundle $F(M)$ relates to the connection 1-form on the base manifold $M$ via the pullback by a section $\sigma:M\to F(M)$:
$\sigma^{*}\omega:=\begin{pmatrix}{\omega_{1}}^{1}&\cdots&{\omega_{n}}^{1}\\ \vdots&\ddots&\vdots\\ {\omega_{1}}^{n}&\cdots&{\omega_{n}}^{n}\end{pmatrix}:T_{p}M\to\mathfrak{o}(n),$ (A.2)
where $\sigma(p)=(p,e_{1},\dots,e_{n})$. If the Stiefel-Whitney classes $w_{1}(M)$ and $w_{2}(M)$ vanish, the frame bundle $F(M)$ reduces to a spin structure $Sp(M)$ with structure group $Spin(n)$ and homomorphism $\rho:Spin(n)\to SO(n)$.
The local connection form $\hat{\omega}$ on $Sp(M)$ is induced by $\omega$ via $\hat{\omega}=(\rho_{*})^{-1}\circ\sigma^{*}\omega$. For the Levi-Civita connection (with ${\omega_{j}}^{i}=-{\omega_{i}}^{j}$), this simplifies to
@@
where $e_{i}e_{j}\in Cl(T_{p}M)$ is the Clifford product. The spinor bundle is built from the $Spin(n)$-representation $\tau:Spin(n)\to GL(\Psi)$. Let $Sp(M)$ be a spin structure on $M$. Using the representation $\tau$, the associated vector bundle $S(M):=Sp(M)\times_{\tau}\Psi$ is defined by quotienting $Sp(M)\times\Psi$ under the equivalence relation $(p\cdot g,v)\sim(p,\tau(g)v)$ for $g\in Spin(n)$. $S(M)$ is a vector bundle with fiber $\Psi$, whose sections are spinor fields.
The covariant derivative of a spinor field $\Psi$ is given by[56]:
$D\Psi=d\Psi+\hat{\omega}\cdot\psi=d\Psi+\frac{1}{4}\sum_{ij}{\omega_{i}}^{j}(e_{i}\cdot e_{j}-e_{j}\cdot e_{i})\cdot\Psi,$ (A.4)
where $\cdot$ denotes the Clifford product. In terms of gamma matrices, this can be rewritten in a more conventional form by defining $\Gamma^{a}:=\eta^{ab}e_{b}\cdot$ in the Lorentzian manifold:
$D\Psi=d\Psi+\frac{1}{4}\sum_{ab}{\omega_{a}}^{b}[\Gamma^{a},\Gamma^{b}]\Psi.$ (A.5)
Here, $\eta={rmdiag\ }(1,-1,-1,-1,-1)$ is the metric signature, and the spinor space inner product is defined as $(\Psi_{1},\Psi_{2}):=i\Psi_{1}^{\dagger}\Gamma^{0}\Psi_{2}$. This formulation satisfies the Clifford algebra relations:
${\Gamma^{a},\Gamma^{b}}=-2\eta^{ab},\quad(\Psi_{1},\Gamma^{a}\Psi_{2})=-(\Gamma^{a}\Psi_{1},\Psi_{2})$ (A.6)
The Dirac operator is given by $\Gamma^{a}D_{e_{a}}$, which leads to the equation of motion for a free fermion of mass $m\in\mathbb{R}$:
@@
In particular, when the metric takes the simple diagonal form
$g=\sum_{a}\eta^{aa}g_{aa}dx^{a}\otimes dx^{a},$ (A.8)
the connection 1-form can be expressed explicitly. First, we define the orthonormal frame and its dual coframe:
$\theta^{a}=\sqrt{g_{aa}}dx^{a},\quad e_{a}=\frac{1}{\sqrt{g_{aa}}}\frac{\partial}{\partial x^{a}}.$ (A.9)
Using Cartan’s structure equation
$d\theta^{a}=\theta^{b}\wedge\omega_{b}{}^{a}$, (A.10)
we obtain the explicit expression for the connection 1-form $\omega_{b}{}^{a}$.
@@
${\omega_{b}}^{a}=\frac{1}{\sqrt{g_{bb}}}\frac{\partial\sqrt{g_{aa}}}{\partial x^{b}}dx^{a}-\frac{1}{\sqrt{g_{aa}}}\frac{\partial\sqrt{g_{bb}}}{\partial x^{a}}dx^{b},~~else.$ (A.11)
@@
$D_{e_{c}}\Psi=\frac{1}{\sqrt{g_{cc}}}\frac{\partial\Psi}{\partial x^{c}}+\frac{1}{4}\sum_{ab}\braket{{\omega_{a}}^{b},e_{c}}[\Gamma^{a},\Gamma^{b}]\Psi.$ (A.12)
@@
In this article, we adopt the abbreviated notation:
$D_{a}\equiv D_{e_{a}}$ (A.14)
for the covariant derivative along the frame field $e_{a}$.
@@ Appendix B Chern class
The Chern class is a cohomology class in ${H_{\text{dR}}}^{2k}(M;\mathbb{R})$ that characterizes the topological structure of a complex vector bundle $(E,M,\pi)$. We begin with the fundamental concept of invariant polynomials.
@@
Let $V$ be a finite-dimensional vector space. For $k\geq 1$, consider a multilinear map:
$f:\underbrace{V\times\cdots\times V}_{k\text{ times}}\to\mathbb{R},\quad(v_{1},\ldots,v_{k})\mapsto f(v_{1},\ldots,v_{k})$ (B.1)
satisfying:
@@
    Symmetry under permutation: $f(v_{\sigma(1)},\ldots,v_{\sigma(k)})=f(v_{1},\ldots,v_{k})$ for all $\sigma\in S_{k}$.
The space of such maps forms a vector space $S^{k}(V^{*})$, with $S^{0}(V^{*}):=\mathbb{R}$.
For a Lie group $G$ with Lie algebra $\mathfrak{g}$, a polynomial $f\in S^{k}(\mathfrak{g}^{*})$ is called invariant if:
$f(\text{Ad}_{g^{-1}}X_{1},\ldots,\text{Ad}_{g^{-1}}X_{k})=f(X_{1},\ldots,X_{k})\quad\forall g\in G$ (B.2)
The space of degree-$k$ invariant polynomials is denoted:
$I^{k}(G):={f\in S^{k}(\mathfrak{g}^{*}):f\text{ is invariant}}$ (B.3)
with $I^{*}(G):=\bigoplus_{k\geq 0}I^{k}(G)$.
For $V$-valued differential forms $\eta_{1},\ldots,\eta_{k}$ on a manifold, expressed in a basis $E_{i}$} as $\eta_{i}=\eta_{i}^{j}E_{j}$, we extend the definition:
$f(\eta_{1},\ldots,\eta_{k}):=f(E_{j_{1}},\ldots,E_{j_{k}})\eta_{1}^{j_{1}}\wedge\cdots\wedge\eta_{k}^{j_{k}}$ (B.4)
For our purposes, the properties of $I^{k}(GL(r,\mathbb{R}))$ are particularly significant.
Proposition B.1 (Differential Properties of Invariant Polynomials)
Let $f\in I^{k}(GL(r,\mathbb{R}))$ be an invariant polynomial. Then:
@@
    For all $X,X_{i}\in\mathfrak{gl}(r,\mathbb{R})$,
    $\sum_{i=1}^{k}f(X_{1},\ldots,[X_{i},X],\ldots,X_{k})=0$ (B.5)
@@
    For $\mathfrak{gl}(r,\mathbb{R})$-valued differential forms $\eta,\eta_{i}$ of odd degree,
    $\sum_{i=1}^{k}f(\eta_{1},\ldots,[\eta_{i},\eta],\ldots,\eta_{k})=0$ (B.6)
@@
The invariance of $f$ implies that for any $t\in\mathbb{R}$ and $X\in\mathfrak{gl}(r,\mathbb{R})$:
$f(e^{-tX}X_{1}e^{tX},\ldots,e^{-tX}X_{k}e^{tX})=f(X_{1},\ldots,X_{k})$ (B.7)
Differentiating at $t=0$ yields the first result. The second property follows by expanding $\eta$ and $\eta_{i}$ in a basis of $\mathfrak{gl}(r,\mathbb{R})$ and applying the first property.
Now consider a vector bundle $(E,M,\pi)$ with structure group $GL(r,\mathbb{R})$, equipped with a connection $D$. Locally, the connection acts on a frame $e_{i}$} of $E$ as:
$De_{i}={\omega_{i}}^{j}\otimes e_{j}$ (B.8)
where $\omega_{j}{}^{i}$} are the connection 1-forms. These can be organized into a matrix-valued 1-form:
@@
The connection form $\omega$ is a $\mathfrak{gl}(r,\mathbb{R})$-valued 1-form, whose curvature $\Omega$ is a $\mathfrak{gl}(r,\mathbb{R})$-valued 2-form defined by:
$\Omega:=d\omega+\frac{1}{2}[\omega,\omega].$ (B.10)
For any invariant polynomial $f\in I^{k}(GL(r,\mathbb{R}))$, we define:
$f(\Omega):=\underbrace{f(\Omega,\ldots,\Omega)}_{k\text{ times}}.$ (B.11)
By the invariance property, $f(\Omega)$ is well-defined globally on the base manifold $M$. This construction has the following crucial properties:
Proposition B.2 (Properties of Characteristic Forms)
For any connection on $(E,M,\pi)$, the form $f(\Omega)$ satisfies:
  1. 1.
    It is closed: $f(\Omega)\in{\Omega_{cl}}^{2k}(M)$
  2. 2.
    Its cohomology class $f(\Omega)]\in{H_{\text{dR}}}^{2k}(M)$ is independent of the connection choice
@@
For closedness, we compute:
$df(\Omega)$ $=\sum_{i=1}^{k}f(\Omega,\ldots,d\Omega,\ldots,\Omega)$
$=\sum_{i=1}^{k}f(\Omega,\ldots,[\Omega,\omega],\ldots,\Omega)\quad\text{(Bianchi identity)}$
$=0\quad\text{(by Proposition 2.2)}.$
For connection independence, consider an interpolation $\omega_{t}=(1-t)\omega_{0}+t\omega_{1}$ between two connections, with $\eta=\omega_{1}-\omega_{0}$. Then:
$f(\Omega_{1})-f(\Omega_{0})$ $=\int_{0}^{1}\frac{d}{dt}f(\Omega_{t})dt$
$=k\int_{0}^{1}f(d\eta+[\eta,\omega_{t}],\Omega_{t},\ldots,\Omega_{t})dt$
$=d\left(k\int_{0}^{1}f(\eta,\Omega_{t},\ldots,\Omega_{t})dt\right)=:d\Phi.$
This construction gives rise to the Chern-Weil homomorphism:
$w(E,\cdot):I^{*}(G)\to{H_{\text{dR}}}^{*}(M),\quad f\mapsto[f(\Omega)].$ (B.12)
Definition B.2 (Chern Classes)
For the unitary group $U(n)\subset GL(n,\mathbb{C})$ with Lie algebra $\mathfrak{u}(n)={A\in\mathfrak{gl}(n,\mathbb{C})|A=-A^{\dagger}}$, consider the characteristic polynomial:
$\det\left(\lambda I_{n}-\frac{1}{2\pi i}A\right)=\sum_{k=0}^{n}c_{k}(A)\lambda^{n-k}.$ (B.13)
The coefficients $c_{k}$ satisfy:
  • Homogeneity: $c_{k}(mA)=m^{k}c_{k}(A)$
  • Ad-invariance: $c_{k}(gAg^{-1})=c_{k}(A)$ for $g\in U(n)$
  • Reality: $c_{k}(A)\in\mathbb{R}$ for $A\in\mathfrak{u}(n)$
The $k$-th Chern class of $E$ is:
$c_{k}(E):=[c_{k}(\Omega)]\in{H_{\text{dR}}}^{2k}(M;\mathbb{R})$ (B.14)
which is a topological invariant independent of the connection.
Appendix C Euler characteristic and Chern number
Following the definition of Chern classes, we now examine a concrete example in 2-dimensional systems where the Chern number directly relates to the Euler characteristic. This connection demonstrates how Chern classes capture fundamental topological properties of physical systems.
@@ Appendix C Euler characteristic and Chern number
[…]
For this $U(1)$-bundle $V_{n}$, the first Chern class takes a particularly simple form. From the characteristic polynomial
$\det\left(\lambda-\frac{A}{2\pi i}\right)=c_{0}(A)\lambda+c_{1}(A),\quad A\in\mathfrak{u}(1),$ (C.3)
we obtain the topological invariant:
$c_{1}(V_{n})=-\frac{[\Omega]}{2\pi i}\in H^{2}_{\text{dR}}(\mathbb{T}^{d}),$ (C.4)
where $\Omega$ is the curvature 2-form.
The natural choice of connection (Berry connection) in this physical context is:
[…]
with corresponding covariant derivative
[…]
The connection we defined satisfies all requirements of a principal $U(1)$-connection. Through Cartan’s structure equation, we obtain the curvature form:
[…]
which leads to the first Chern class for normalized eigenvectors:
[…]
For 2D systems ($d=2$), we define the Chern number of the energy band $E_{n}$ as:
[…]
This integral is well-defined since exact forms integrate to zero by Stokes’ theorem.
The eigenspace
[…]
admits a natural realification: